OpenStudy (anonymous):

PLEASE HELP! Medal & Fan Given. Martina combined different amounts of water, vinegar, and rubbing alcohol to make 38 milliliters (ml) of a solution for a science experiment. The amount of water was x ml. The amount of vinegar was 8 ml less than that of water, and the amount of rubbing alcohol was half the amount of vinegar. This is represented in the equation below: x + (x - 8) + one half (x - 8) = 38 What was the difference in the amounts of water and rubbing alcohol Martina combined? A. 8 ml B. 14 ml C. 18 ml D. 20 ml

OpenStudy (calculusxy):

To first solve this problem, you need to figure out the value of x (water).

OpenStudy (calculusxy):

Here, I will make it easier to see the equation: \[x +(x - 8) + \frac{ 1 }{ 2 }(x-8) = 38\]

OpenStudy (calculusxy):

Do you know how to solve for x?

OpenStudy (anonymous):

Just plug in the values in the choices?

OpenStudy (calculusxy):

Well no, really. These are the differences between the amount of water and the amount of vinegar.

OpenStudy (calculusxy):

Try to simplify the equation. Meaning, combine the like terms together.

OpenStudy (anonymous):

Is the answer 20 ml because I plugged in that for x and it equaled 38

OpenStudy (calculusxy):

Wait. Give me a moment because I need to solve the problem myself to check whether or not you are correct.

OpenStudy (calculusxy):

Okay I think that I got an answer. But we will answer this together to make sure that you get it. Since you said that 20ml, plugged in to x, equals to 38, we will say that 20ml will be the amount of water that you have. Understood?

OpenStudy (calculusxy):

To figure out the amount of rubbing alcohol, we will have to first do x - 8 (which will result us with the amount of vinegar).

OpenStudy (anonymous):

i know i did all of the other choices and they didn't equal 38

OpenStudy (calculusxy):

Okay. So how much will the amount of vinegar equal to?

OpenStudy (calculusxy):

Remember that the problem says, " The amount of vinegar was 8 ml less than that of water..."

OpenStudy (calculusxy):

Are you here? I need your cooperation..

OpenStudy (anonymous):

sorry, i'm kinda busy that's why i'm replying slowly.

OpenStudy (calculusxy):

Anyways, since we have 20 ml to figure out the amount of vinegar, we need to do 20-8=12. Half of 12 equals to 6. Therefore ,the amount of rubbing alcohol is 6ml. The difference is therefore 20-6=14..

OpenStudy (anonymous):

Wait... so what's the answer?

OpenStudy (calculusxy):

14ml (i believe)

OpenStudy (anonymous):

okay, thank you :)

OpenStudy (calculusxy):

welcome :)